Question 1

Suppose you made $65,000 last year for coming in second place in a beauty contest. During this time the consumer price index (CPI) increased from 109 to 115.

a) What is the inflation rate?

b) Does this signify inflation, deflation, disinflation, or hyperinflation? Why?

c) If the second-place prize for the beauty contest has an escalator clause built-in, how much should the prize now be?
Answer #1

Inflation is percentage change in CPI = (115-109)*100/109 = 5.50%

This is an increase in price therefore inflation.

If there is an escalator clause, then the price is adjusted for inflation.

When CPI was 109, prize was 65000. If now CPI is 115, Prize is 65000*115/109 = 68577.98